# thanks: The National Key Research and Development Program of China (Project No. 2022YFA1602001)

### v1

@@ VI Conclusion
In this study, we provided a detailed introduction to the unique design and superior performance of the module, which holds meaningful value for the process design of plastic scintillator detectors with WLS-fibers. Additionally, we proposed a batch quality inspection process and established a set of standards for evaluating the performance of plastic scintillator modules, offering valuable experience and reference for quality inspection in other related experiments. The specific summary is as follows: The more the muon hits the position towards the two ends of the PS strip, the higher the total light yield but with a stronger asymmetry. When the muon hits the center of the PS strip (minimum effective light yield region along the strip, air coupling between SiPM and PS), for a 2000mm-PS module, the most probable signal strength output from one end is at least greater than 40.8 p.e. For a 1500mm-PS module, it is at least greater than 51.5 p.e. All the modules meet the requirement proposed by JUNO-TAO that the light yield to a muon should be greater than 40 p.e. With optical grease to couple SiPM and WLS-fiber, it can increase the effective light yield by 12.5%.
The muon detection efficiency is checked with different thresholds on the output of single end, sum or coincidence of two ends of the module. It can serve as a universal indicator for muon detection, and we measured and calculated in detail the detection efficiency with muon by nine positions along the length direction of the PS module. Overall, the 1500-mm PS detection efficiency is higher than that of 2000-mm PS. All PS module detection efficiencies are above 99.3% and meet TAO’s requirement of detection efficiency exceeding 99%.

### v2

@@ VI Conclusion
In this study, we provided a detailed introduction to the unique design and performance of the module, which holds meaningful value for the process design of plastic scintillator detectors with WLS-fibers. Additionally, we proposed a batch quality inspection process and established a set of standards for evaluating the performance of plastic scintillator modules, offering valuable experience and reference for quality inspection in other related experiments. The specific summary is as follows: The more the muon hits the position towards the two ends of the PS strip, the higher the total light yield but with a stronger asymmetry. When the muon hits the center of the PS strip (minimum effective light yield region along the strip, air coupling between SiPM and PS), for a 2000mm-PS module, the most probable signal strength output from one end is at least greater than 40.8 p.e. For a 1500mm-PS module, it is at least greater than 51.5 p.e. All the modules meet the requirement proposed by JUNO-TAO that the light yield to a muon should be greater than 40 p.e. With optical grease to couple SiPM and WLS-fiber, it can increase the effective light yield by 12.5%.
The muon detection efficiency is checked with different thresholds on the output of single end, sum or coincidence of two ends of the module. Three types of detection efficiency have been defined by us, comprehensively evaluating the performance of the plastic scintillator. Overall, the 1500-mm PS detection efficiency is better than that of 2000-mm PS. In "AND" mode, when the threshold is set at 20 photoelectrons, the detection efficiency exceeds 99.3%, and at 15 photoelectrons, it surpasses 99.6%, with clear differentiation between background and muon signals. These features and advantages result in our plastic scintillator performing exceptionally well in experiments, providing a reliable detection solution for other experiments.
VII Acknowledgements
This work is supported by the General Program of National Natural Science Foundation of China (Project No.12075087), the National Key Research and Development Program of China (Project No.2022YFA1602002), the Strategic Priority Research Program of the Chinese Academy of Sciences (Project No. XDA10011102). Guang Luo also greatly appreciates the support from Guangdong Provincial Key Laboratory of Advanced Particle Detection Technology.